# Insights from Large-Scale Communications: Initiation

### v1

@@ 2.2 Information of large-scale Structures
It is imperative to consider that the mass distribution of the universe has been discretely characterized based solely on the presence or absence of galaxies. Since galaxies occupy a minuscule portion of the total space, it is more probable to encounter empty regions than substantial overdensities. When comparing a galaxy cluster to an individual void, a unit volume extracted from a void contains less information than a unit volume from a cluster, owing to the vast size of voids. Nonetheless, the collective contribution from voids proves significant due to their higher likelihood of occurrence. Regardless of whether the unit volume is extracted from an under-dense or overdense region, uncertainty is always lost. Hence, departures from the mean background would lead to a change of information, which goes into shaping newly formed structures. However, the extent and direction of this information loss vary between overdense and under-dense structures.
At this point, we make another crucial assertion by imposing the holographic principle on the cube in consideration and define the total weighted   Information Gain   for any $i^{th}$ voxel as
${\cal G}_{i}(\delta)$ $=$ $\frac{A_{c}}{4l_{P}^{2}}\left[\,\frac{p_{i}(\delta)\left\{\,I_{i}(\delta)-\bar{I}\,\right\}}{\bar{p}\bar{I}\,U_{0}}\,\right].$ (12)
Here $A_{c}=6\left(\frac{L}{1+z}\right)^{2}$ is the surface area of the cube with its center at redshift $z$. The Planck length $l_{P}$ is derived from gravitational constant ($\,G\,$), Planck constant ($\,\hbar\,$) and speed of light ($\,c\,$) as $l_{P}=\sqrt{\frac{G\hbar}{c^{3}}}$. The total excess probabilities within the cube is denoted as $U_{0}=\sum_{i}\exp\{\bar{I}-I_{i}\}$. For a completely homogeneous distribution, the information gain for any voxel and the entire cube would be zero. We rewrite Equation 12 in terms of excess information $\Delta I_{i}=(I_{i}-\bar{I})$,
${\cal G}_{i}(\delta)$ $=$ $\frac{K}{U_{0}\bar{I}}\left[\,\Delta I_{i}\,\exp\{\,-\Delta I_{i}\,\}\,\right].$ (13)
$K=\frac{3u_{I}L^{2}}{2(1+z)^{2}l_{P}^{2}}$ is a constant for a specific cube, positioned at a given redshift. The factor $u_{I}=\frac{2^{-43}}{\ln 2}$ is used to convert the results from nats to terabytes. For overdense structures, we get +ve ${\cal G}$ and -ve ${\cal G}$ is obtained for under-dense spaces. As we have stated before, the growth of structures forces a system to lose information, which subsequently gets stored as the configuration of the emergent structure.
The right panels of Figure 2 shows ${\cal G}$ as a function of $\delta$ in symmetric log scale, for different choices of $\gamma$ and ${\bar{f}_{\cal M}}$. from the bottom right panel of this figure it is clear that ${\cal G}$ is not sensitive to ${\bar{f}_{\cal M}}$. For an infinitesimal ${\bar{f}_{\cal M}}$, Equation 10 and Equation 11 leads to
$I(\,\delta\,)=\bar{I}+\frac{\delta}{\delta_{c}},$ (14)
@@
with $\,\delta_{c}=\left[\gamma\ln\left(1+\frac{1}{\gamma}\right)\right]^{-1}$ Equation 13 becomes
${\cal G}(\delta)=\frac{K}{U_{0}\bar{I}}\left[\,\frac{\delta}{\delta_{c}}\exp\{-\frac{\delta}{\delta_{c}}\}\,\right],$ (15)
@@
The function ${\cal G}(\delta)$ has four characteristic points min, null, max and fall, at $\delta=-1$, $\delta=0$, $\delta=\delta_{c}$ and $\delta\sim 10\delta_{c}$ respectively. These critical points which define the curve are marked in the bottom right panel of Figure 2. Now, $\delta_{c}$ remains a free parameter unless we constrain it through observational evidence. In this work, we choose $\gamma={8}\times 10^{-4}$ to set $\delta_{c}\sim 175$ as this critical value represents the density threshold required for the virialization of a system. Past this threshold, the system configuration which has gradually developed over time through the externally acquired information, begins to be overridden by the internal stochasticity. Further increment in density contrast would cause the information gain to be suppressed and finally, it will diminish at the core ($\delta\to\infty$).
2.3 Information density and spectrum of information
@@
We conduct a numerical analysis to search the scale of homogeneity in the SDSS galaxy distribution. After a cube of size $L$ is carved out from the SDSS survey volume ( as described in subsection 3.1 ), it is divided into $N_{s}^{3}$ voxels of size $\Delta x$. The density contrast for each voxel is then found from the effective galaxy count. We estimate information gain for each voxel using the relation in Equation 15. In this section, we do not impose $\bar{f}_{\cal M}\to 0$ and choose a finite value of $\bar{f}_{\cal M}$, as we are trying to validate the Cosmological principle itself. Figure 3 shows Information Gain at different locations within one of the cubes in Sample 1. It also presents how information gain correlates with galaxy clustering and respective probabilities. We use $N_{s}=50$ grids, setting $\bar{f}_{\cal M}={8}\times 10^{-6}$. So, there are $100$ virtual chunks in use.
@@
After we have the Information gain calculated at each of the $N_{v}$ grids, we find the dimensionless 3D Information spectrum as the Fourier transform of information gain in the unitary form
$\cal S(\mathbf{k})$ $=$ $\int_{\mathbb{R}^{3}}{\Sigma(\mathbf{x})\,\exp\left\{-2\pi i\,\mathbf{k}\cdot\mathbf{x}\right\}}\,\,d^{3}x.$ (16)
Here $\Sigma(\,\mathbf{x}_{i}\,)=\frac{{\cal G}_{SNC}(\,\mathbf{x}_{i}\,)}{\Delta x^{3}}$ is the information density with ${\cal G}_{SNC}$ as the shot noise corrected information gain ( see A ). $\cal S$ is a dimensionless complex quantity, computed using the Discrete Fourier Transform (DFT), i.e.
${\cal S}_{[l,m,n]}=\frac{1}{\sqrt{N_{v}}}\sum_{l=0}^{N_{s}-1}\sum_{m=0}^{N_{s}-1}\sum_{n=0}^{N_{s}-1}\Sigma_{[a,b,c]}\;\exp\left\{\frac{-2\pi i}{N_{s}}(al+bm+cn)\right\}$ (17)
@@
$k_{m}$ $=$ $k_{N}\left[\frac{2m}{N_{s}}-1\right];\quad\quad\text{for $\left\{m=0,1,2,\ldots,(N_{s}-1)\right\}$}$ (18)
@@
Now, our goal is to find $\cal S$ as a function of $k$. We proceed by considering the entire available 3D $k$-space to be comprised of $N_{b}$ spherical shells centred at ($\frac{N_{s}}{2}$,$\frac{N_{s}}{2}$,$\frac{N_{s}}{2}$). Any $j^{th}$ shell is extended from $\frac{(j-1)\,\,k_{N}}{N_{b}}$ to $\frac{j\,\,k_{N}}{N_{b}}$, if $j$ takes values from $1$ to $N_{b}$. On the other hand, for any of the grid [$l$,$m$,$n$] in Fourier space, the magnitude of $k$ is found as $|\,k\,|=\sqrt{\,k_{l}^{2}+k_{m}^{2}+k_{n}^{2}\,}$. The spherical bin that the grid belongs to is then $b=\lfloor\frac{N_{b}\,|\,k\,|}{k_{N}}\rfloor$. For each bin, we calculate the 1D Information spectrum, $\bar{\cal S}(k)$ by taking a bin-wise average over $\cal S$, where each spherical bin has a specific value of $k$, associated with it. $\bar{\cal S}(k)$ now quantifies the average information gain across different frequencies.
2.4 Information sharing across different length scales
We perform an inverse-Fourier transform on $\bar{\cal S}$, normalizing it by the Nyquist frequency, to measure the Information gain shared by the space on different length scales,
@@ 2.4 Information sharing across different length scales
$W(k)=\frac{i_{0}\left(\,\,\beta\sqrt{1-\left[\frac{k}{k_{N}}\right]^{2}}\,\,\right)}{i_{0}\left(\,\,\beta\,\,\right)}.$ (20)
@@
Here, $i_{0}$ is the modified Bessel function of the zeroth-order and $\beta=14$. The simplified form of Equation 19 in terms of DFT would be
${{{\cal G}^{M}}}_{[a]}$ $=$ $\frac{1}{\sqrt{N_{b}}}\,\,\left[\sum_{m=0}^{N_{s}-1}W_{[m]}\,\,\bar{\cal S}_{[m]}\;\exp\left\{\frac{2\pi i\,a\,m}{N_{b}}\right\}\right]$ (21)
@@
Mutual gain essentially tells how much surplus information produced from the LSS formation is mutually shared by any two points in space, on average. The magnitude of ${{\cal G}^{M}}(r)$ signifies the extent of uncertainty reduction attributed to one point in space, due to the information stored at another point separated by a distance $r$. In the context of large-scale statistical homogeneity, by definition, the cause of information gain itself becomes redundant. Consequently, one would expect the mutual gain to exhibit a diminishing trend across the scale of homogeneity.
This section is concentrated on validating the existence of cosmic homogeneity and determining the length scale associated with it. So, the idea of a homogeneous universe has not been adopted thus far. To maintain causal consistency, our approach initially involved checking for the existence of cosmic homogeneity before its further application in the next part of the analysis.
2.5 Evolution of global information
$D_{+}(\,a\,)=\frac{5\Omega_{m0}E(a)}{2}\int^{a}_{0}{\frac{da^{\prime}}{{a^{\prime}}^{3}E(a^{\prime})^{3}}},$ (22)
with $E(a)=\frac{H(a)}{H_{0}}$ as the scaled Hubble parameter.
Let us now break down the assertion of $\bar{f}_{\cal M}\to 0$ which leads to Equation 15. Keeping $\gamma$ at a finite value, one has to keep adding voxels up to a point where all the virtual chunks that span the Universe are taken care of. This would require an enormously large number of voxels or an infinitesimal $\bar{f}_{\cal M}$. Hence, $\bar{f}_{\cal M}\to 0$ automatically imposes the assumption of cosmic homogeneity in this case.
At this point we introduce the global information entropy of the Universe as
${\cal I}^{U}=\frac{K^{U}}{\bar{I}}\left[\frac{\int_{\mathbb{R}^{3}}\,p(\,\delta\,)\,I(\,\delta\,)\,d^{3}x}{\int_{\mathbb{R}^{3}}\,p(\,\delta\,)\,d^{3}x}\right],$ (23)
with the redshift-dependent global maxima for information given as
$K^{U}(z)=\frac{u_{I}c^{5}}{4G\hbar H_{0}^{2}}\left[\Omega_{m0}(1+z)^{3}+\Omega_{\Lambda}\right]^{-1}.$ (24)
Now, $K^{U}(0)$ is the maximum amount of information that the present universe with a homogeneous matter distribution would contain, and this is $\sim 2.81\times 10^{147}\,\,terabytes$ of information. The global information entropy measures the effective information content in the Universe at a given time, considering the total information gain ( or loss ) caused by structure formation. Note that ${\cal G}^{U}$ being marginalized over space, is a function of the scalefactor (or redshift) only.
Before going forward, let us also introduce the functional exponentially weighted moment (EWM) at this point. The $k^{th}$ order EWM is defined as
${\cal A}_{k}[\,\eta\,]=\frac{\int_{\mathbb{R}^{3}}\,\{\,\eta(\mathbf{x},a)\,\}^{k}\,\exp\{\,-\eta(\mathbf{x},a)\,\}d^{3}x}{\int_{\mathbb{R}^{3}}\,\exp\{\,-\eta(\mathbf{x},a)\,\}d^{3}x}.$ (25)
With $a$ as the scalefactor and $\eta(\mathbf{x},a)=\frac{\delta(\mathbf{x},a)}{\delta_{c}}$ as the scaled contrast, we can now use Equation 14 and Equation 15 and rewrite Equation 23 as
[…]
$=$ $K^{U}+{\cal G}^{U}\,\,,$
where $K^{U}$ is the upper information bound and ${\cal G}^{U}(\,a\,)$ is the global information gain. We also term $(-\,{\cal G}^{U})$ as the global information loss hereafter. The rate of change of global information entropy with the scalefactor is
$\frac{d{\cal I}^{U}}{da}$ $=$ $\frac{dK^{U}}{da}+\left(\frac{{\cal A}_{1}}{\bar{I}}\frac{dK^{U}}{da}+\frac{K^{U}}{\bar{I}}\frac{d{\cal A}_{1}}{da}\right).$ (27)
The term inside the parenthesis on the right side of this equation gives the rate of change of information gain, i.e. $\frac{d{\cal G}^{U}}{da}$. Now, let us denote $\frac{d\ln K^{U}}{d\ln a}=\frac{3\Omega_{m}}{\Omega_{m}+\Omega_{\Lambda}}={\cal E}(\Omega_{m})$ with $\Omega_{m}=\Omega_{m0}a^{-3}$ and $\frac{1}{2}\left(\frac{5a}{3D_{+}}-1\right)=\Gamma(a)$. This allows the logarithmic growth rate ($\frac{d\ln D_{+}}{d\ln a}$) to be expressed as $\,\,f(\Omega_{m})=\Gamma(a)\,{\cal E}(\Omega_{m})\,\,$, and leads to
$\frac{d{\cal G}^{U}}{da}$ $=$ $\frac{K^{U}\,{\cal A}_{1}\,{\cal E}}{a\,\bar{I}}\left[1+\Gamma\left\{1+{\cal A}_{1}-\frac{{\cal A}_{2}}{{\cal A}_{1}}\right\}\right],$ (28)
through the $1^{st}$ derivative of the $1^{st}$ EWM ( see B ).
Equation 28 can be restructured in the following forms
$\frac{d\ln{\cal G}^{U}}{d\ln a}$ $=$ ${\cal E}(\Omega_{m})\left[1+\Gamma\left\{1+{\cal A}_{1}-\frac{{\cal A}_{2}}{{\cal A}_{1}}\right\}\right],$ (29)
$\frac{d\ln{\cal G}^{U}}{d\ln H}$ $=$ $-2\left[1+\Gamma\left\{1+{\cal A}_{1}-\frac{{\cal A}_{2}}{{\cal A}_{1}}\right\}\right],$ (30)
$\frac{d\ln{\cal G}^{U}}{d\ln D_{+}}$ $=$ $\frac{1}{\Gamma}+\left\{1+{\cal A}_{1}-\frac{{\cal A}_{2}}{{\cal A}_{1}}\right\}$ (31)
Note that Equation 30 presents the identity $-\frac{1}{(1+q)}\frac{d\ln{\cal G}^{U}}{d\ln a}$ with $q=-\frac{a\ddot{a}}{\dot{a}^{2}}$ being the deceleration parameter. The quantities listed above in combination are useful for validating the $\Lambda CDM$ model and cosmological parameter estimation since $A_{1}$ and $A_{2}$ can be readily determined from observations. However, this work only focuses on the overall evolution of information gain for a log-normal density distribution.
Let us now go back to Equation 25 and perform Taylor series expansion of the exponential terms, which gives the $k^{th}$ EWM in terms of the moments of $\eta(a)$; i.e.
${\cal A}_{k}=\frac{\sum_{i=0}^{\infty}{\frac{(\,-1\,)^{i}}{i!}}\,\langle\,\eta(a)^{i+k}\,\rangle}{\sum_{i=0}^{\infty}{\frac{(\,-1\,)^{i}}{i!}}\,\langle\,\eta(a)^{i}\,\rangle},$ (32)
where $\langle\eta(a)\rangle=0$, and $\langle\eta(a)^{2}\rangle=\frac{D_{+}^{2}(a)\sigma_{8}^{2}}{\delta_{c}^{2}}$.
Now, if we consider ($1+\delta$) to have a log-normal distribution with the skewness and kurtosis characterized by the $S_{p}$ parameters (bernardeau95), then the $3^{rd}$ and $4^{th}$ order moments of $\eta$ will be found as
$\langle\eta(a)^{3}\rangle$ $=$ $\frac{S_{3}\sigma^{4}}{\delta_{c}^{3}},$ (33)
$\langle\eta(a)^{4}\rangle$ $=$ $\frac{S_{4}\sigma^{6}+3\sigma^{4}}{\delta_{c}^{4}},$ (34)
where $\sigma\,=\,D_{+}\sigma_{8}$. The $S_{p}$ parameters at any given redshift $z$, are found as $S_{3}(z)=S_{3,0}(1+z)^{\alpha_{3}}$ and $S_{4}(z)=S_{4,0}(1+z)^{\alpha_{4}}$. The present values of $S_{3}$ and $S_{4}$ are $S_{3,0}=3+\sigma_{8}^{2}$ and $S_{4,0}=16+15\sigma_{8}^{2}+6\sigma_{8}^{4}+\sigma_{8}^{6}$ respectively. $\alpha_{3}$ and $\alpha_{4}$ control the rate of change of the skewness and kurtosis of the distribution at a given redshift. In this work, we have used $\alpha_{3}=-1.1$ and $\alpha_{4}=-2.3$ (einasto21) to study the evolution of global information entropy from $z=20$ to $z=0$.

### v2

@@ 2.2 Information associated with density perturbations
It is imperative to consider that the mass distribution of the universe has been discretely characterized based solely on the presence or absence of galaxies. Since galaxies occupy a minuscule portion of the total space, it is more probable to encounter empty regions than substantial overdensities. When comparing a galaxy cluster to an individual void, owing to the abundance of voids in the Universe, a unit volume extracted from a void contains less information than its counterpart from a cluster. However, the collective contribution of information from voids is significant due to the higher likelihood. The departures from the mean background would lead to a change in information, which in the case of overdense structures goes into its internal configuration changing the degrees of freedom of the constituents. On the other side, the expanding space would offer more states for the chunks to be arranged, resulting in a growth of uncertainty and information. So, the extent and direction of the change in information entropy would vary between overdense and under-dense structures.
2.3 Information entropy and large-scale structure formation
At this point, we define the normalized  Information Entropy for the entire region as
${\cal E}^{I}$ $=$ $\frac{1}{\bar{I}}\,\,\biggl{[}\frac{\int_{V}\,\,d^{3}x\,\,\,p(\delta(\mathbf{x}))\,\,\,I(\delta(\mathbf{x}))\,}{\,\,\int_{V}\,\,d^{3}x\,\,\,p(\delta(\mathbf{x}))}\biggr{]}.$ (2.11)
However, the effective change in information entropy is measured through the  Entropic gain, defined as
${\cal G}^{I}$ $=$ $\frac{1}{\bar{I}}\,\,\biggl{[}\frac{\int_{V}\,\,d^{3}x\,\,\,\Delta I(\delta)\,\,\exp\{-\Delta I(\delta)\}\,}{\,\,\int_{V}\,\,d^{3}x\,\,\,\exp\{-\Delta I(\delta)\}}\biggr{]};$ (2.12)
Note that ${\cal E}^{I}=1+{\cal G}^{I}$, implying a reduction in total information entropy for a negative entropic gain. Assessment of the entropic gain does not require any presumption on the probability distribution of $\delta$. Despite having no prior knowledge regarding the profile of $\delta$, the collective excess information can be estimated by applying the appropriate weights to each elementary volume element based on the associated density contrast.
For a cubic region discretely divided into a finite number of grids with spacing $\Delta x$, one can measure The specific information gain $\Phi\,(\,\mathbf{x}_{i}\,)$, i.e. the excess amount of information (weighted) per unit volume around $\mathbf{x}_{i}$. With $U_{0}=\sum_{i}\,\exp\{\,-\Delta I\}$, the specific information gain can be expressed as
$\Phi\,(\,\mathbf{x}_{i}\,)\,$ $=$ $\frac{1}{\bar{I}\,U_{0}\,\Delta x^{3}}\biggl{[}\,\,\,\Delta I\,(\delta\,(\mathbf{x}_{i})\,)\,\,\exp\{\,-\Delta I\,(\,\delta(\mathbf{x}_{i})\,)\,\}\,\,\,\biggr{]}.$ (2.13)
For a completely homogeneous distribution, the specific information gain at any given point as well as the total entropic gain of the entire volume would be zero. We obtain +ve and -ve $\Phi$ for underdense and overdense structures respectively. The right panels of Figure 2 show $\bar{I}\,U_{0}\Delta x^{3}\Phi$ or $\Delta Ie^{-\Delta I}$ as a function of $\delta$ in symmetric log scale, whereas the left and middle panels show the probability and information associated to different overdensities. The Top panels show the dependence on density contrast when $\gamma$ is varied. The bottom panels show the dependence on $\bar{f}_{\cal M}$. The function $\Delta Ie^{-\Delta I}$ has four characteristic points min, null, max and fall defining the curve, which are marked in the bottom right panel of Figure 2. The min, null and max points are located at $\delta=-1$, $\delta=0$, $\delta=\left[\gamma\ln\left(1+\frac{1}{\gamma}\right)\right]^{-1}$ respectively. From the bottom right panel, it is clear that $\Phi$ is not sensitive to the number of volume elements when $\gamma$ is fixed. To incorporate the effect of structure formation all over the Universe, one can keep $\gamma$ fixed and start adding voxels in all directions to extend the cube’s volume to accommodate all the virtual chunks that span the Universe. This would require an enormously large number of voxels or an infinitesimal $\bar{f}_{\cal M}$. Imposing the condition ${\bar{f}_{\cal M}}\to 0$ leads to
$I(\,\delta\,)=\bar{I}+\delta\left[\gamma\ln\left(1+\frac{1}{\gamma}\right)\right].$ (2.14)
@@ 2.3 Information entropy and large-scale structure formation
With $\,\delta_{c}=\left[\gamma\ln\left(1+\frac{1}{\gamma}\right)\right]^{-1}$ and $\eta(\mathbf{x})=\frac{\delta(\mathbf{x})}{\delta_{c}}$ Equation 2.12 becomes
${\cal G}^{U}$ $=$ $\frac{1}{\bar{I}}\,\,\biggl{[}\frac{\int_{{\mathbb{R}}^{3}}\,\,d^{3}x\,\,\,\eta(\mathbf{x})\,\,\exp\{-\eta(\mathbf{x})\}\,}{\,\,\int_{{\mathbb{R}}^{3}}\,\,d^{3}x\,\,\,\exp\{-\eta(\mathbf{x})\}}\biggr{]};$ (2.15)
@@
Here ${\cal G}^{U}$ is the global information gain. $\delta_{c}$ remains a free parameter unless it is constrained through experimental evidence. In this work, we choose $\gamma={8}\times 10^{-4}$ to set $\delta_{c}\sim 175$ as this critical value represents the density threshold required for the virialization of a system. Past this threshold, the system configuration which has gradually developed over time through the externally acquired information, begins to be overridden by the internal stochasticity. Further increment in density contrast would cause the specific information gain to be suppressed and finally, it will diminish at the core ($\delta\to\infty$).
2.4 Evolution of the global information entropy
$D_{+}(\,a\,)=\frac{5\,\Omega_{m}(a)\,a^{3}\,\,E(a)^{3}}{2}\int^{a}_{0}{\frac{da^{\prime}}{{a^{\prime}}^{3}E(a^{\prime})^{3}}},$ (2.16)
where $\Omega_{m}(a)=\frac{\Omega_{m0}\,a^{-3}}{E(a)^{2}}$ is the matter density parameter at the scale factor $a$ with the normalised Hubble parameter as $E(a)=\frac{H(a)}{H_{0}}=\sqrt{\Omega_{m0}\,a^{-3}+(1-\Omega_{m0})}$.
Before going forward, let us introduce the functional exponentially weighted moment (EWM). The $k^{th}$ order EWM is defined as
${\cal A}_{k}[\,\eta\,]=\frac{\int_{\mathbb{R}^{3}}\,\{\,\eta(\mathbf{x},a)\,\}^{k}\,\exp\{\,-\eta(\mathbf{x},a)\,\}d^{3}x}{\int_{\mathbb{R}^{3}}\,\exp\{\,-\eta(\mathbf{x},a)\,\}d^{3}x}.$ (2.17)
With $a$ as the scalefactor and $\eta(\mathbf{x},a)=\frac{\delta(\mathbf{x},a)}{\delta_{c}}$ as the scaled contrast, we can now find the global entropic gain (Equation 2.15) in terms of the first EWM, i.e
${\cal G}^{U}=\frac{{\cal A}_{1}[\,\eta\,]}{\bar{I}}$ (2.18)
Taking the derivative w.r.t the scale factor ( see subsection A.1 ) leads to
$\frac{d{\cal G}^{U}}{da}$ $=$ $\frac{f(\Omega_{m})}{a\,\bar{I}}\left[{\cal A}_{1}^{2}+{\cal A}_{1}-{\cal A}_{2}\right],$ (2.19)
where $f(\Omega_{m})$ is the logarithmic growth rate $\frac{d\ln D_{+}}{d\ln a}$. Equation 2.19 can also be restructured to find the following identities
$\Gamma_{D+}=\frac{d\ln D_{+}}{d\ln{\cal G}^{U}}$ $=$ $\left[1+{\cal A}_{1}-\frac{{\cal A}_{2}}{{\cal A}_{1}}\right]^{-1},$ (2.20)
$\Gamma_{a}=\frac{d\ln a}{d\ln{\cal G}^{U}}$ $=$ $\frac{1}{f(\Omega_{m})}\left[1+{\cal A}_{1}-\frac{{\cal A}_{2}}{{\cal A}_{1}}\right]^{-1},$ (2.21)
$\Gamma_{H}=\frac{d\ln H}{d\ln{\cal G}^{U}}$ $=$ $-\frac{1+q}{f(\Omega_{m})}\left[1+{\cal A}_{1}-\frac{{\cal A}_{2}}{{\cal A}_{1}}\right]^{-1}$ (2.22)
Note that $q$ in Equation 2.21 is the decelearation parameter, i.e. $q=-\frac{a\ddot{a}}{\dot{a}^{2}}$. At a given instant, the logarithmic slope $\Gamma_{X}=\frac{d\ln X}{d\ln{\cal G}^{U}}$ for an arbitrary variable $X$ relates the change in $X$ to the change in entropic gain through $\Delta X\,=\,\Gamma_{X}\,\Delta{\cal G}^{U}\,\left(\,\frac{X\,}{{\cal G}^{U}}\,\right)$. And for a constant $\Gamma_{X}$ we also have $X=X_{0}\left(\,{\cal G}^{U}\,\right)^{\Gamma_{X}}$. Here the factor $\Gamma_{X}$ regulates the growth of $X$ at any instant, by controlling the sensitivity of $X$ on ${\cal G}^{U}$.
Let us now go back to Equation 2.17 and perform Taylor series expansion of the exponential terms, which gives the $k^{th}$ EWM in terms of the moments of $\eta(a)$; i.e.
${\cal A}_{k}=\frac{\sum_{i=0}^{\infty}{\frac{(\,-1\,)^{i}}{i!}}\,\langle\,\eta(a)^{i+k}\,\rangle}{\sum_{i=0}^{\infty}{\frac{(\,-1\,)^{i}}{i!}}\,\langle\,\eta(a)^{i}\,\rangle},$ (2.23)
where
[…]
We consider ($1+\delta$) to have a log-normal distribution with the skewness and kurtosis characterised by the $S_{p}$ parameters [77]. Hence, the $3^{rd}$ and $4^{th}$ order moments of $\eta$ will be
[…]
Here $\sigma\,=\,D_{+}\sigma_{8}$ and the $S_{p}$ parameters at any given redshift $z$, are found as $S_{3}(z)=S_{3,0}(1+z)^{\alpha_{3}}$ and $S_{4}(z)=S_{4,0}(1+z)^{\alpha_{4}}$; where the present values of $S_{3}$ and $S_{4}$ are given as $S_{3,0}=3+\sigma_{8}^{2}$ and $S_{4,0}=16+15\sigma_{8}^{2}+6\sigma_{8}^{4}+\sigma_{8}^{6}$ respectively. $\alpha_{3}$ and $\alpha_{4}$ control the rate of change of the skewness and kurtosis of the distribution at a given redshift. In this work, we have used $\alpha_{3}=-1.1$ and $\alpha_{4}=-2.3$ [78] to study the evolution of global information entropy from $z=20$ to $z=0$.
The quantities $\Gamma_{D_{+}}$, $\Gamma_{a}$ and $\Gamma_{H}$ presented in Equation 2.20, Equation 2.21 and Equation 2.22 in combination are useful for validating the $\Lambda CDM$ model and cosmological parameter estimation since ${\cal A}_{1}$ and ${\cal A}_{2}$ can be readily determined from observations. However, this work only focuses on the overall evolution of entropic gain for a log-normal density distribution through the $\Gamma$ parameters connecting the entropic gain with the expansion rate of the Universe and the growth rate of the LSS.
2.5 Spatial distribution of Information
@@
We conduct a numerical analysis to search the scale of homogeneity in the SDSS galaxy distribution. After a cube of size $L$ is carved out from the SDSS survey volume ( as described in subsection 3.1 ), it is divided into $N_{s}^{3}$ voxels of size $\Delta x$. The density contrast for each voxel is then found from the effective galaxy count. We estimate the specific information gain for each voxel using the relation in Equation 2.13. In this section, we do not impose $\bar{f}_{\cal M}\to 0$ and choose a finite value of $\bar{f}_{\cal M}$, as we are trying to validate the Cosmological principle itself. $\bar{f}_{\cal M}$ indirectly takes the cosmic homogeneity and isotropy into account, to assert that $\gamma$ remains constant when the boundary is extended to the horizon. that Figure 3 shows the specific information gain at different locations within one of the cubes in Sample 1. It also presents how entropic gain correlates with galaxy clustering and respective probabilities. We use $N_{s}=50$ grids, setting $\bar{f}_{\cal M}={8}\times 10^{-6}$. So, there are $100$ virtual chunks in use.
2.5.1 Information spectrum
@@
After we have found $\Phi_{SNC}(\,\mathbf{x}_{i}\,)$, i.e. the shot noise corrected specific information gain ( see subsection A.2 ) at each of the $N_{v}$ grids, we find the dimensionless 3D Information spectrum as its Fourier transform
$\cal S(\mathbf{k})$ $=$ $\int_{\mathbb{R}^{3}}{\Phi_{SNC}(\mathbf{x})\,\exp\left\{-2\pi i\,\mathbf{k}\cdot\mathbf{x}\right\}}\,\,d^{3}x.$ (2.26)
$\cal S$ is a dimensionless complex quantity that quantifies the amplitude of the entropic gain in different frequency ranges. It is numerically evaluated using the Discrete Fourier Transform (DFT), i.e.
${\cal S}_{[l,m,n]}=\frac{1}{\sqrt{N_{v}}}\sum_{l=0}^{N_{s}-1}\sum_{m=0}^{N_{s}-1}\sum_{n=0}^{N_{s}-1}\Phi_{[a,b,c]}\;\exp\left\{\frac{-2\pi i}{N_{s}}(al+bm+cn)\right\}$ (2.27)
@@
$k_{m}$ $=$ $k_{N}\left[\frac{2m}{N_{s}}-1\right];\quad\quad\text{for $\left\{m=0,1,2,\ldots,(N_{s}-1)\right\}$}$ (2.28)
@@
Now, our goal is to find $\cal S$ as a function of $k$. We proceed by considering the entire available 3D $k$-space to be comprised of $N_{b}$ spherical shells centred at ($\frac{N_{s}}{2}$,$\frac{N_{s}}{2}$,$\frac{N_{s}}{2}$). Any $j^{th}$ shell is extended from $\frac{(j-1)\,\,k_{N}}{N_{b}}$ to $\frac{j\,\,k_{N}}{N_{b}}$, if $j$ takes values from $1$ to $N_{b}$. On the other hand, for any of the grid [$l$,$m$,$n$] in Fourier space, the magnitude of $k$ is found as $|\,k\,|=\sqrt{\,k_{l}^{2}+k_{m}^{2}+k_{n}^{2}\,}$. The spherical bin that the grid belongs to is then $b=\lfloor\frac{N_{b}\,|\,k\,|}{k_{N}}\rfloor$. For each bin, we calculate the spherically averaged Information spectrum, $\bar{\cal S}(k)$ by taking a bin-wise average over $\cal S$, where each spherical bin has a specific value of $k$, associated with it. $\bar{\cal S}(k)$ now quantifies the average entropic gain across different frequencies.
2.5.2 Information sharing across different length scales
We perform an inverse-Fourier transform on $\bar{\cal S}$, normalizing it by the Nyquist frequency, to measure the entropic gain shared by the space on different length scales,
@@ 2.5.2 Information sharing across different length scales
$W(k)=\frac{i_{0}\left(\,\,\beta\sqrt{1-\left[\frac{k}{k_{N}}\right]^{2}}\,\,\right)}{i_{0}\left(\,\,\beta\,\,\right)}.$ (2.30)
@@
Here, $i_{0}$ is the modified Bessel function of the zeroth-order and $\beta=14$. The simplified form of Equation 2.29 in terms of DFT would be
${{\cal G}^{M}}_{[a]}$ $=$ $\frac{1}{\sqrt{N_{b}}}\,\,\left[\sum_{m=0}^{N_{s}-1}W_{[m]}\,\,\bar{\cal S}_{[m]}\;\exp\left\{\frac{2\pi i\,a\,m}{N_{b}}\right\}\right]$ (2.31)
@@
Mutual gain essentially tells how much of the extra information ( on average ) produced from the LSS formation is mutually shared by any two points in space. The magnitude of ${{\cal G}^{M}}(r)$ signifies the degree of uncertainty reduction attributed to one point in space, due to the exposure of the information stored at another point separated by a distance $r$. In the context of large-scale statistical homogeneity, by definition, the entropic gain diminishes with the absence of perturbations in the density field. Consequently, one would expect the mutual gain to exhibit a diminishing trend across the scale of homogeneity.